OpenStudy (anonymous):

HELP PLEASE! PLEASE! PLEASE! I WILL GIVE YOU A MEDAL Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

PART 1 : if x is the number of strawberries and y is the number of Grapes : 2 times the strawberries you have and the number of strawberries you have would add and make 10 fruits. Also, three times the strawberries you have and 4 times the grapes you have would add and make 25 fruits.

OpenStudy (anonymous):

good example

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

PART 2 : 2x+y=10......1 3x+4y=25.....2 from equation 1 x=10-y/2 put it in 2nd equ 3(10-y/2)+4y=25 30-3y/2+4=25 now by LCM 30-3y+8y/2=25 by cross multiply 30+5y=50 5y=20 Y = 4 now put y=4 in equ 1 to get value of x 2x+4=10 2x=10-4 2x=6 X = 3 ((: I Hope This Was Hopeful